OpenStudy (anonymous):

Find an equation of the circle whose diameter has endpoints (3,-5)and (1,1)?

OpenStudy (anonymous):

Do you know what the general equation of a circle is?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

(x-h)^2+(y-k)^2=r^2

OpenStudy (shadowlegendx):

I remember this, but so vaguely...

OpenStudy (anonymous):

@chachi123 (x-h)^2+(y-k)^2=r^2 is right, your answer is correct

OpenStudy (anonymous):

fint the mid point of those two end points, which will become your centre value. this centre value has co-ordinates (h,k)